OpenStudy (anonymous):

A monkey of mass 20kg is holding a vertical rope.The rope can break when a mass of 25kg is suspended from it.What is the maximum acceleration with which monkey can climb up along the rope?

OpenStudy (anonymous):

assuming the value of g is 10ms^-2, the maximum weight that the rope can hold is \[25kg \times10 {Nkg^-1}= 250 N\]. (though it should be less than 250, I'm not so sure about the maimum part) using F=ma, \[a= {250N \over 20 kg} =12.5 ms^-2\] The maximum acceleration of the monkey is 12.5ms^-2

OpenStudy (anonymous):

I think, if the monkey doesn't climb but only hang on a rope, the pulling strength upon it would be 200N (if g equals 10). If the monkey starts climb up it will produce additional pulling on the rope, so the total pulling will be m(g+a). By the data it can't be above 250N (the rope will break off then). Thus, a = 250/20 - 10 = 2,5 (m/s^2).